The weight of 1000 children average 50 kg and the standard deviation is 5kg. What is the Area under the normal curve that corresponds to the z-score of 55


kg?

1
Expert's answer
2022-04-26T13:16:29-0400

Let "X" be a random variable denotes the children's weight: "X\sim N(\mu, \sigma^2)"


"Z=\dfrac{X-\mu}{\sigma}\sim N(0, 1)"

"z=55"


"P(Z=55)=0"

"P(Z<55)=P(Z\le 55)\approx1"

"P(Z>55)=P(Z\ge 55)\approx0"